# ConVerSum: A Contrastive Learning based Approach for Data-Scarce Solution of Cross-Lingual Summarization Beyond Direct Equivalents

### v1

@@ 5.3 ConVerSum vs LLMs
Based on our experiments on LLM for low-resource languages demonstrated in Table 9, we have some observations on LLMs. These are as follows:
5.3.1 Gemini 1.5 Pro:
@@ 5.3.3 ChatGPT 4o:
When reviewing the capabilities of the latest and most powerful LLM, ChatGPT-4o, which was launched on May 13, 2024, it is stated that, while this model is capable of handling low-resource languages, it still has limits. It cannot handle large documents (more than 700 words) for several low-resource languages, such as Burmese, emphasizing areas where further improvements are required.
5.4 LaSE vs BERTScore
Instances, where ConVerSum’s LaSE is higher than the baseline, highlight its strength in cross-lingual summary generation preserving the overall structure with meaning across different languages. Moreover, higher BERTScore focuses on higher semantic similarity. This dual-metric evaluation delivers significant insights to satisfy the demands of different application scenarios.
5.5 Overall Discussion
@@
ConVerSum performs better in both high- and low-resource scenarios, and large- and small-scale datasets, and consistently beats the baseline models. It even performs better than GPT-3.5 and almost as well as GPT-4o in low-resource language settings, outperforming the capabilities of current LLMs. The observation that ConVerSum performs well even with limited training data for low-resource languages highlights the efficacy of the contrastive learning approach used. These outcomes validate the robustness and flexibility of ConVerSum, highlighting the ability of contrastive learning to improve model performance even in resource-constrained scenarios.
6 Conclusion

### v2

@@ 4.12. Discussion on LLM for CLS
Based on our experiments on LLM for low-resource languages, we have some observations on LLMs. These are as follows:
4.12.1. Gemini 1.5 Pro:
@@ 4.12.3. ChatGPT 4o by Prompting Approach:
When reviewing the capabilities of the latest and most powerful LLM, ChatGPT-4o, which was launched on May 2024, it is stated that, while this model is capable of handling low-resource languages, it still has limits. It cannot handle large documents (more than 700 words) for several low-resource languages, such as Burmese, emphasizing areas where further improvements are required.
4.12.4. ChatGPT 4o Using API:
GPT 4o outperforms for several language pairs specifically it shows significantly better performance for high-resource languages. However, GPT-4o’s performance is somewhat weaker for language pairings with limited resources where it involves idioms or syntax of those languages.
5. Result Analysis in Different Aspects
This section analyzes our findings based on different aspects like low and high-resource scenarios, and large and small-scale dataset scenarios. Furthermore, the detailed analysis based on the evaluation metrics is discussed.
5.1. For Training on High-Resource Large-Scale Dataset
5.1.1. Comparison between mT5 and Flan-T5 based Model
For all cases, the Flan T5 model as a baseline and the proposed model using Flan T5 show significantly lower performance than ConVerSum using mT5 and also for the mT5 itself as a baseline. The robustness of ConVerSum heavily relies upon the quality of candidate generation used in contrastive learning. As there is no Flan-T5 checkpoint specifically fine-tuned on cross-lingual pairs, the generated candidates fail to show diverse quality and languages for Flan-T5. Though Flan-T5 is multilingual, it sometimes shows biases in high-resource languages like English.
5.1.2. ConVerSum using mT5
While training on the high-resource large-scale dataset as discussed in Section 4.6, ConVerSum outperforms the baseline in LaSE for most source-target language pairs, except Bengali-English and Tigrinya-English, where it performs similarly to the baseline. For BERTScore, our model’s performance is nearly identical to the baseline mT5 for a few cases and others outperform the baseline.
5.2. For Training on Low-Resource Small-Scale Dataset
While training on the low-resource small-scale dataset as discussed in Section 4.8, ConVerSum outperforms the baseline for most of the language pairs, and only for a few instances, it shows almost identical performance to baseline.
5.3. Comparison between LaSE and BERTScore for CLS
In almost all cases, ConVerSum outperforms the baseline for LaSE rather than BERTScore except in a few cases. For CLS evaluation, LaSE is generally considered a better choice than BERTScore as LaSE prioritizes meaning and fluency across languages, which is more important than just word-level similarity. BERTScore only focuses on higher semantic similarity. There are some results where BERTScore improves and LaSE doesn’t. This scenario indicates the meaning similarity is improved but is unable to guarantee the summary length and target language. For example, we want to generate a summary for an English source document to target Bengali. We get two summaries as mentioned in Table 19.
Table 19. Comparison between LaSE and BERTScore for CLS
No Summary Analysis
1
Word overlap results in a high BERTScore,
although Bengali and English are mixed together.
Lower LaSE for language confidence.
2
Lower BERTScore due to lack of direct
word overlap but provides a correct
CLS higher LaSE
From Table 19, Summary 1 still receives a high BERTScore though there is a mixing of source and target languages. The LaSE score is lower in this case because it ensures summaries are generated in the target language. Summary 2 receives a lower BERTScore as there is a lack of direct word overlap though it provides a correct cross-lingual summary. Therefore, LaSE receives a high score in this case maintaining its criteria for CLS evaluation (Bhattacharjee et al., 2023). This dual-metric evaluation delivers essential insights to satisfy the demands of different application scenarios.
5.4. Overall Discussion
@@
ConVerSum performs better in both high- and low-resource scenarios, and large- and small-scale datasets, and consistently beats the baseline models. It even performs better than GPT-3.5 and almost as well as GPT-4o in low-resource language settings, outperforming the capabilities of current LLMs. The observation that our model performs well even with limited training data for low-resource languages highlights the efficacy of the contrastive learning approach used. These outcomes validate the robustness and flexibility of our model, highlighting the ability of contrastive learning to improve model performance even in resource-constrained scenarios.
6. Conclusion
@@
Acknowledgements.
This work was funded by the Research and Innovation Centre for Science and Engineering (RISE), BUET.